Duncan hits a volleyball out of bonds on a path given by the equation
h= -o.1d^2+d+1. The height of the volleyball is h metres, and the horizontal distance from Duncan is d metres. What is the maximum height reached by the ball?

Answer by stanbon(48558)

h= -0.1d^2+d+1. The height of the volleyball is h metres, and the horizontal distance from Duncan is d metres. What is the maximum height reached by the ball?
--------------
Max occurs when d = -b/2a = -1/(2*-0.1) = -1/-0.2 = 5
----------------
Max height = h(d) = -0.1d^2 + d + 1
h(5) = -0.1(5)^2 + 5 + 1
h(5) = 3.5
=================
Cheers,
Stan H.